OpenStudy (anonymous):

What fraction would show 3.4 as the quotient of two integers?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

34/10 or 17/5

OpenStudy (anonymous):

help

OpenStudy (anonymous):

what do you mean 34/10 @eliassaab

OpenStudy (anonymous):

anyone please help

OpenStudy (anonymous):

Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x – y ≥ 7 x + 2y < 5 A. (6, 1) B. (8, –1) C. (6, 2) D. (6, –2)

OpenStudy (anonymous):

please help

OpenStudy (mathmale):

@poptartmn : Jean initiated this dialogue by posting a question. Her question, not yours, is the focus here. Please post your own questions separately. Thank you. mathmale (Moderator)

OpenStudy (mathmale):

Please post your questions as new questions, now.

OpenStudy (anonymous):

I'm a girl but thank you

OpenStudy (anonymous):

i mean guy

OpenStudy (anonymous):

ahahaha oops

OpenStudy (mathmale):

Jean, if you're still interested: 3.4 is a mixed number (part integer: 3, and part decimal fraction: 0.4. Our goal is to convert that to a ratio of two integers. You could write 3.4 as \[\frac{ 3.4 }{ 1 }\] and then multiply both numerator and denominator by 10. What happens? Please respond.

OpenStudy (anonymous):

I'm not good at math @mathmale

OpenStudy (mathmale):

Can you simplify this?\[\frac{ 3.4 }{ 1 }*\frac{ 10 }{ 10 }\]

OpenStudy (mathmale):

Hints: multiply 3.4 by 10. Multiply that 1 in the denominator by 10. What do you get (in the form of a fraction which is the ratio of two integers?

OpenStudy (isaiah.feynman):